A poll found that 53​% of a sample of 1170 teens in a certain large country go online several times a day.​ (Treat this sample as a simple random​ sample.) Find the margin of error for this poll if we want 90​% confidence in our estimate of the percent of teens in the country who go online several times a day. [ ]? - Round to one decimal place as​ needed
Answer #1

Giren, n=1170 ² = 0.53 , a=0.10 margin of error = 20.05* | P(1-P) = 1644854 * (0:53) (1-0-53) 1170 = 1.644854 * 0.014591298 =